OpenStudy (anonymous):

helllpppp!? change each decimal to a fraction? 1.) 0.7 2.) 0.21 3.) 0.03 4.) 0.1 5.) 0.119

OpenStudy (anonymous):

Lol Alekxos helping every problem

OpenStudy (anonymous):

1. 7/10 2. 21/100 3. 3/100 4. 1/10 5. 119/1000 Each decimal place represents a fraction with a denominator of 10 for the first place to the right of the decimal, 100 for the second place, 1,000 for the third place, and so on.

OpenStudy (anonymous):

1) 7/10 Write the decimal number as a fraction (over 1) 0.7 = 0.7 / 1 Multiplying by 1 to eliminate 1 decimal places we multiply top and bottom by 1 10's Numerator (N) N = 0.7 × 10 = 7 Denominator (D) D = 1 × 10 = 10 N / D = 7 / 10 Simplifying our fraction = 7/10

OpenStudy (anonymous):

Yeah, I'm trying to :D

OpenStudy (anonymous):

2) Write the decimal number as a fraction (over 1) 0.21 = 0.21 / 1 Multiplying by 1 to eliminate 2 decimal places we multiply top and bottom by 2 10's Numerator (N) N = 0.21 × 10 × 10 = 21 Denominator (D) D = 1 × 10 × 10 = 100 N / D = 21 / 100 Simplifying our fraction = 21/100

OpenStudy (anonymous):

Write the decimal number as a fraction (over 1) 0.03 = 0.03 / 1 Multiplying by 1 to eliminate 2 decimal places we multiply top and bottom by 2 10's Numerator (N) N = 0.03 × 10 × 10 = 3 Denominator (D) D = 1 × 10 × 10 = 100 N / D = 3 / 100 Simplifying our fraction = 3/100

OpenStudy (anonymous):

Write the decimal number as a fraction (over 1) 0.1 = 0.1 / 1 Multiplying by 1 to eliminate 1 decimal places we multiply top and bottom by 1 10's Numerator (N) N = 0.1 × 10 = 1 Denominator (D) D = 1 × 10 = 10 N / D = 1 / 10 Simplifying our fraction = 1/10

OpenStudy (anonymous):

Write the decimal number as a fraction (over 1) 0.119 = 0.119 / 1 Multiplying by 1 to eliminate 3 decimal places we multiply top and bottom by 3 10's Numerator (N) N = 0.119 × 10 × 10 × 10 = 119 Denominator (D) D = 1 × 10 × 10 × 10 = 1000 N / D = 119 / 1000 Simplifying our fraction = 119/1000

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

kinda! thank you!

OpenStudy (anonymous):

No problem :)